The velocity of a moving object is given by the function

𝒅𝑽=𝟓𝒕^2+2t-4/t^2

a) Integrate the function and find the distance travelled in metres between t=2 and t=5

b) Calculate the distance travelled using a numerical method and compare your answer with part A.


1
Expert's answer
2022-01-11T12:49:04-0500

a)


"V=5t^2+2t-\dfrac{4}{t^2}"

"distance=\displaystyle\int_{2}^{5}(5t^2+2t-\dfrac{4}{t^2})dt"

"=\big[\dfrac{5t^3}{3}+t^2+\dfrac{4}{t}\big]\begin{matrix}
 5 \\
 2
\end{matrix}"

"=\dfrac{5(5)^3}{3}+(5)^2+\dfrac{4}{5}-(\dfrac{5(2)^3}{3}+(2)^2+\dfrac{4}{2})"

"=214.8 (units)"

b) The Simpson's 1/3 rule

Approximate the integral "\displaystyle\int_{2}^{5}(5t^2+2t-\dfrac{4}{t^2})dt" with "n=6"  using the Simpson's rule.


"\displaystyle\int_{a}^{b}V(t)dt=\dfrac{\Delta t}{3}(V(t_0)+4V(t_1)+2V(t_2)+..."

"+2V(t_{n-2})+4V(t_{n-1})+V(t_n))"

We have that  "V(t)=5t^2+2t-\dfrac{4}{t^2}, a=2, b=5, n=6"

Then "\Delta t=\dfrac{b-a}{n}=\dfrac{5-2}{6}=0.5"

Divide the interval "[2,5]" into "n=6" subintervals of the length "\Delta t=0.5" with the following endpoints: "a=2, 2.5, 3, 3.5, 4, 4.5, 5=b."

Now, just evaluate the function at these endpoints.


"V(t_0)=V(2)=23"

"4V(t_1)=4V(2.5)=142.44"

"2V(t_2)=2V(3)=\dfrac{910}{9}"

"4V(t_3)=4V(3.5)=\dfrac{13313}{49}"

"2V(t_4)=2V(4)=175.5"

"4V(t_5)=4V(4.5)=\dfrac{35657}{81}"

"V(t_6)=V(5)=134.84"

"\dfrac{1}{6}(23+142.44+\dfrac{910}{9}+\dfrac{13313}{49}+175.5"

"+\dfrac{35657}{81}+134.84)\approx 214.7991(units)"

Theanswer for distance from part A "214.8" units and from part B "214.799" units are very close


"\dfrac{|214.799-214.8|}{214.8}\times 100\%=0.0005\%"